CONTINUOUS-WAVE CAVITY RINGDOWN STUDY OF THE ^{14}N_{2}^{+} MEINEL SYSTEM 2-1 BAND AND THE FIRST POSITIVE BAND SYSTEM OF N_{2}^{∗}

\underline{SUSANNA L. WIDICUS WEAVER}, Departments of Chemistry and Astronomy, University of Illinois at Urbana-Champaign, Urbana, IL 61801; BOGDAN NEGRU, MICHAEL WICZER, and BRIAN A. TOM, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801; JOSHUA P. DiGANGI, Department of Chemistry, University of Wisconsin, Madison, Wisconsin 53706; BENJAMIN J. McCALL, Departments of Chemistry and Astronomy, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

We have observed the 2-1 vibronic band of the N_{2}^{+} A^{2}Π_{u}-X^{2}Σ_{g}^{+} system using continuous-wave cavity ringdown spectroscopy (cw-CRDS). The N_{2}^{+} was formed in a positive column discharge of a mixture of N_{2} and argon, and spectra were collected between 937 and 984 nm with a tunable external cavity diode laser. A total of 112 N_{2}^{+} lines have been observed. In addition, more than 400 lines from the first positive band system of N_{2}^{∗} were also observed during this experiment. The relative intensities of the N_{2}^{∗} and N_{2}^{+} bands were found to change with discharge cell pressure, and so each spectral region was observed at two different pressures to aid in distinguishing the spectra of these species. Both the N_{2}^{+} and N_{2}^{∗} spectra have been assigned and the molecular constants have been determined for each species, and the interpretation of these spectra will be reported.